OpenStudy (anonymous):

Let sin A = 12/13 with 90º≤A≤180º and tan B = -4/3 with 270º≤B≤360º. Find tan (A + B). I need help PLEASE

OpenStudy (anonymous):

I need the fraction

OpenStudy (anonymous):

? PLEASE!

myininaya (myininaya):

\[\tan(A+B)=\frac{\sin(A+B)}{\cos(A+B)}=\frac{\sin(A)\cos(B)+\sin(B)\cos(A)}{\cos(A)\cos(B)-\sin(A)\sin(B)}\] So we need to know the following: sin(A);sin(B);cos(A);cos(B)